0 14.P010. is 158 dB a) Calculate the intensity of the sound wave generated by the jet at the given location. W/m2
Answer #1

a)
we know,

Sound Intensity level, beta = 10*log(I/Io) dB

158 = 10*log(I/Io)

15.8 = log(I/Io)

10^15.8 = I/Io

I = 10^15.8*Io

I = 10^15.8*10^-12

I = 6309 W/m^2 <<<<-----------------Answer

b) I/I_threshold = 6309/100

= 63

the intensity of jet plane = 63 times of Intebsity of threshold pain